# Data-Adaptive Graph Framelets with Generalized Vanishing Moments for Graph Signal Processing

### v2

@@ II-C Construction of $\mathcal{G}$-Framelet Systems
In summary, the nodes of $\mathcal{G}^{\prime}$ represent clusters and the edge weight on $\mathcal{G}^{\prime}$ is determine by summing of the edge weight among nodes of two clusters. We called $\mathcal{G}^{\prime}$ a coarse-grained graph of $\mathcal{G}$ [44]. Based on the coarse-grained graph, we can give the definition of multi-graph partition trees.
Definition 3.
A multi-graph partition tree $\mathscr{T}_{J}({\mathcal{G}})$ of $\mathcal{G}=(\mathcal{V},\mathcal{E},\bm{W})$ with $J+1,J\in\mathbb{N}$ levels is a partition tree $\mathscr{T}_{J}({\mathcal{V}})$ as defined in Definition 1 such that each $j\in\{0,\dots,J\}$ is associated with a coarse-grained graph $\mathcal{G}^{j}=(\mathcal{V}^{j},\mathcal{E}^{j})$ of $\mathcal{G}$ and $\mathcal{V}^{j}=\{\mathcal{S}_{j,k}\,|\,k\in[n_{j}]\}$. In particular, $\mathcal{G}^{J}\equiv\mathcal{G}$, where we consider a vertex in $\mathcal{G}^{J}$ as a cluster of singleton.
An intuitive and equivalent interpretation of the definition above is that a multi-graph partition tree is a partition tree by generated successively forming coarsed-grained graphs $\mathcal{G}^{J-1},\mathcal{G}^{J-1}\dots,\mathcal{G}^{0}$ such that $\mathcal{G}^{j}$ is a coarse-grained graph of $\mathcal{G}^{j+1},j\in[J-1]$. Since $\mathscr{T}_{J}(\mathcal{G})$ is associated with a partition tree $\mathscr{T}_{J}(\mathcal{V})$, we can construct $\mathcal{V}$-framelet systems as in Definition 2. We have the following definition.
Definition 4.
Let $\mathscr{T}_{J}(\mathcal{G})$ be a multi-graph partition tree associated with a graph $\mathcal{G}=(\mathcal{V},\mathcal{E},\bm{W})$. Let {$r_{j}$, $\bm{A}^{[j,k]}$, $\bm{B}^{[j,k]}\,|\,j=0,\ldots,J-1;k\in[n_{j}]\}$ be the set of integers and filter matrices satisfying the assumptions of Theorem 1. Then for $J_{0}=0,\ldots,J-1$, the $\mathcal{V}$-framelet system $\mathcal{F}_{J_{0}}^{J}(\mathcal{V})$ as in (14) is a tight frame for $\mathbb{V}_{J}=\mathrm{span}\{\bm{e}_{k}\,|\,k\in[n]\}=L_{2}(\mathcal{G})$, which we call a (graph-involved) $\mathcal{G}$-framelet (GIF) system. In such a case, we simply denote $\mathcal{F}_{J_{0}}^{J}(\mathcal{G}):=\mathcal{F}_{J_{0}}^{J}(\mathcal{V})$ to indicate the role played by $\mathcal{G}$. In particular, we define for the special case $J_{0}=0$, the GIF system $\mathcal{F}_{0}^{J}(\mathcal{G})$ as
$\mathsf{GIF}_{J}(\mathcal{G}):=\mathcal{F}_{0}^{J}(\mathcal{G})=\bm{\Phi}_{0}\oplus\bm{\Psi}_{0}\oplus\cdots\oplus\bm{\Psi}_{J-1}.$ (18)
Remark 7.
@@
III Graph Framelet Transforms
@@
III-A Graph Framelet Transforms
@@
$\bm{c}^{[j,k]}:=\langle\bm{f},\bm{\Phi}^{[j,k]}\rangle,\quad\bm{d}^{[j,k]}:=\langle\bm{f},\bm{\Psi}^{[j,k]}\rangle,$ (20)
@@
$\left\{\begin{matrix}\bm{c}^{[j,k]}&=\langle\bm{\llparenthesis}\bm{c}^{[j+1,k^{\prime}]}\bm{\rrparenthesis}_{k^{\prime}\in\mathcal{C}_{j,k}},\mathcal{A}^{[j,k]}\rangle,\\ \\ \bm{d}^{[j,k]}&=\langle\bm{\llparenthesis}\bm{c}^{[j+1,k^{\prime}]}\bm{\rrparenthesis}_{k^{\prime}\in\mathcal{C}_{j,k}},\mathcal{B}^{[j,k]}\rangle,\\ \end{matrix}\right.$ (21)
@@
$(\bm{c}^{[j+1,k^{\prime}]})_{k^{\prime}\in\mathcal{C}_{j,k}}=\bm{c}^{[j,k]}\mathcal{A}^{[j,k]}+\bm{d}^{[j,k]}\mathcal{B}^{[j,k]},\$ (22)
@@
On the other hand, in view of the condition $(\mathcal{A}^{[j,k]})^{\top}\mathcal{A}^{[j,k]}+(\mathcal{B}^{[j,k]})^{\top}\mathcal{B}^{[j,k]}=\bm{I}$, together with (21), it is easy to see that (22) holds. ∎
The one-level decomposition and reconstruction relations in Corollary 2 can be applied successively to decompose or reconstruct a graph signal $\bm{f}$. Such multi-level decomposition (forward transform) and reconstruction (backward transform) procedures are summarized in Algorithms 3 and 4, and illustrated in Figures 2 and 3, respectively. Here, in Algorithm 4, the notion $\text{ID}(k^{\prime})$ is defined by sorting $k^{\prime}\in\mathcal{C}_{j,k}$ increasingly to index each $k^{\prime}\in\mathcal{C}_{j,k}$ by a unique integer $\text{ID}(k^{\prime})$ in $\{1,\dots,|\mathcal{C}_{j,k}|\}$ such that $\text{ID}(k^{\prime})<\text{ID}(k^{\prime\prime})\Leftrightarrow k^{\prime}\leq k^{\prime\prime}$.
@@
Figure 2: An illustration of computing coefficient vectors on a non-leaf node of $\mathscr{T}_{J}(\mathcal{G})$. The whole transform is computed iteratively in a bottom-up manner at each non-leaf node of $\mathscr{T}_{J}(\mathcal{G})$
Figure 2: An illustration of computing coefficient vectors on a non-leaf node of $\mathscr{T}_{J}(\mathcal{G})$. The whole transform is computed iteratively in a bottom-up manner at each non-leaf node of $\mathscr{T}_{J}(\mathcal{G})$
Figure 3: An illustration of reconstruction on a non-leaf node of $\mathscr{T}_{J}(\mathcal{G})$. The whole transform is computed iteratively in a top-down manner at each non-leaf node of $\mathscr{T}_{J}(\mathcal{G})$. Here, $[\cdot]_{i,:}$ denotes the operation of extracting the $i$-th row of the matrix.
III-B Computational Complexity
@@
Theorem 2.
[…]
$t\leq|\mathcal{C}_{j,k}|\leq s,\,s\leq t^{\alpha},\,r_{j}\leq r\leq\frac{s}{2},\,m_{j,k}\leq m$ (23)
@@
$+$ $\mathcal{O}(s\cdot r^{J-j})$ (Matrix multiplication with $(\text{{Matrix multiplication with }}\bm{A}^{[j,k]}),$
@@
As for Algorithm 4, it is obvious that lines 4, 5, and 6 together become the reverse of lines 5 and 6 of Algorithm 3, and that the amounts of computation are the same. Thus, the computational complexity of Algorithm 4 is the same as that of Algorithm 3. ∎
Remark 8.
@@
IV Data-adpative Graph Framelet Systems
@@
For determining the low-pass filters, we aim at learning low-pass filters so that the space $\mathbb{V}_{0}=\mathrm{span}(\bm{\Phi}_{0})$ provides the best fitting of the samples in the $l^{2}$ sense. That is, for any signal $\bm{f}\in\mathfrak{F}$, most of the energy of $\bm{f}$ is contained in its projection on $\mathbb{V}_{0}$. In such a way, we could capture the low-dimensional structure of the family $\mathfrak{F}$. Once we obtained the low-pass filters, in view of the generalized vanishing moments, i.e., orthogonality between $\mathbb{V}_{0}$ and $\mathbb{W}_{j}$, most of the coefficients in $\mathbb{W}_{j}$ are expected to be small or even zero. Therefore, in such a case, the $\mathcal{G}$-framelet system $\mathsf{GIF}_{J}(\mathcal{G})$ will provide sparse representations for all signals in $\mathfrak{F}$.
The collection of matrices in $\mathbb{R}^{n\times p}$ such that $\bm{M}^{\top}\bm{M}=\bm{I}$ forms a Riemannian manifold endowed with the Riemannian metric induced from the Euclidean metric of $\mathbb{R}^{n\times p}$. These manifolds are called Stiefel manifolds [45]. We denote such a manifold in $\mathbb{R}^{n\times p}$ by $\mathcal{M}^{[n,p]}$. We next provide the learning of the low- and high-pass filters through solving optimization problems on Stiefel manifolds.
IV-A Learning the Low-pass Filters
@@
More specifically, assuming $r_{j},j=0,\ldots,J-1$ are specified such that $r_{j}\in[\min(|\mathcal{C}_{j,1}|,\dots,|\mathcal{C}_{j,n_{j}})|-1]$, we solve the following optimization problem for the low-pass filters:
$\underset{{\left\{\bm{A}^{[j,k]}\in\mathcal{M}^{[r_{j},|\mathcal{C}_{j,k}|]}\,|\,k\in[n_{j}]\right\}_{j=0}^{J-1}}}{\arg\max}\|\bm{c}^{[0,1]}_{1}\|^{2}+\dots+\|\bm{c}^{[0,1]}_{n_{f}}\|^{2},\$ (24)
@@
where $\bm{c}^{[0,1]}_{i}=\langle\bm{f}_{i},\bm{\Phi}_{0}\rangle$, $i\in[n_{f}]$, are the coefficients with respect to $\bm{\Phi}_{0}$ for $\bm{f}_{i}$, $i\in[n_{f}]$, computed as in Algorithm 3 and $\|\cdot\|_{2}$ is the $l_{2}$ norm. Note that these coefficients only depend on the filters $\bm{A}^{[j,k]}$. By solving the optimization problem in (24), the scaling functions $\bm{\Phi}_{0}$ and the space $\mathbb{V}_{0}$ are simultaneously determined. Moreover, the objective function in (24) is smooth on the product manifold $\mathcal{M}_{\bm{A}}:=\otimes_{j,k}\mathcal{M}^{[r_{j},|\mathcal{C}_{j,k}|]}$ formed by all $\mathcal{M}^{[r_{j},|\mathcal{C}_{j,k}|]}$. Thus, the problem above can be regarded as an optimization problem on Riemannian manifolds and we can choose any Riemannian optimization algorithm that is eligible to apply, e.g., see [45].
IV-B Learning the High-pass Filters

### v3

@@ 2.3 Construction of $\mathcal{G}$-Framelet Systems
In summary, the nodes of $\mathcal{G}^{\prime}$ represent clusters and the edge weight on $\mathcal{G}^{\prime}$ is determined by summing of the edge weight among nodes of two clusters. We called $\mathcal{G}^{\prime}$ a coarse-grained graph of $\mathcal{G}$ [73]. Based on the coarse-grained graph, we can give the definition of multi-graph partition trees.
Definition 2.15.
A multi-graph partition tree $\mathscr{T}_{J}({\mathcal{G}})$ of $\mathcal{G}=(\mathcal{V},\mathcal{E},\bm{W})$ with $J+1,J\in\mathbb{N}$ levels is a partition tree $\mathscr{T}_{J}({\mathcal{V}})$ as defined in Definition 2.6 such that each $j\in\{0,\dots,J\}$ is associated with a coarse-grained graph $\mathcal{G}^{j}=(\mathcal{V}^{j},\mathcal{E}^{j})$ of $\mathcal{G}$ and $\mathcal{V}^{j}=\{\mathcal{S}_{j,k}\,|\,k\in[n_{j}]\}$. In particular, $\mathcal{G}^{J}\equiv\mathcal{G}$, where we consider a vertex in $\mathcal{G}^{J}$ as a cluster of singleton.
An intuitive and equivalent interpretation of the definition above is that a multi-graph partition tree is a partition tree by generated successively forming coarse-grained graphs $\mathcal{G}^{J-1},\mathcal{G}^{J-1}\dots,\mathcal{G}^{0}$ such that $\mathcal{G}^{j}$ is a coarse-grained graph of $\mathcal{G}^{j+1},j\in[J-1]$. Since $\mathscr{T}_{J}(\mathcal{G})$ is associated with a partition tree $\mathscr{T}_{J}(\mathcal{V})$, we can construct $\mathcal{V}$-framelet systems as in Definition 2.7. We have the following definition.
Definition 2.16.
Let $\mathscr{T}_{J}(\mathcal{G})$ be a multi-graph partition tree associated with a graph $\mathcal{G}=(\mathcal{V},\mathcal{E},\bm{W})$. Let $r_{j}$, $\bm{A}^{[j,k]}$, $\bm{B}^{[j,k]}\,|\,j=0,\ldots,J-1;k\in[n_{j}]\}$ be the set of integers and filter matrices satisfying the assumptions of Theorem 2.8. Then for $J_{0}=0,\ldots,J-1$, the $\mathcal{V}$-framelet system $\mathcal{F}_{J_{0}}^{J}(\mathcal{V})$ as in (13) is a tight frame for $\mathbb{V}_{J}=\mathrm{span}\{\bm{e}_{k}\,|\,k\in[n]\}=L_{2}(\mathcal{G})$, which we call a (graph-involved) $\mathcal{G}$-framelet (GIF) system. In such a case, we simply denote $\mathcal{F}_{J_{0}}^{J}(\mathcal{G}):=\mathcal{F}_{J_{0}}^{J}(\mathcal{V})$ to indicate the role played by $\mathcal{G}$. In particular, we define for the special case $J_{0}=0$, the GIF system $\mathcal{F}_{0}^{J}(\mathcal{G})$ as
(17) $\mathsf{GIF}_{J}(\mathcal{G}):=\mathcal{F}_{0}^{J}(\mathcal{G})=\{\bm{\Phi}_{0},\bm{\Psi}_{0},\cdots,\bm{\Psi}_{J-1}\}.$
Remark 2.17.
@@
3 Graph Framelet Transforms
@@
3.1 Graph Framelet Transforms
@@
(19) $\bm{c}^{[j,k]}:=\langle\bm{f},\bm{\Phi}^{[j,k]}\rangle,\quad\bm{d}^{[j,k]}:=\langle\bm{f},\bm{\Psi}^{[j,k]}\rangle,$
@@
[…]
@@
(21) $(\bm{c}^{[j+1,k^{\prime}]})_{k^{\prime}\in\mathcal{C}_{j,k}}=\bm{c}^{[j,k]}\mathcal{A}^{[j,k]}+\bm{d}^{[j,k]}\mathcal{B}^{[j,k]},\\$
@@
On the other hand, in view of the condition $(\mathcal{A}^{[j,k]})^{\top}\mathcal{A}^{[j,k]}+(\mathcal{B}^{[j,k]})^{\top}\mathcal{B}^{[j,k]}=\bm{I}$, together with (20), it is easy to see that (21) holds.
The one-level decomposition and reconstruction relations in Corollary 3.1 can be applied successively to decompose or reconstruct a graph signal $\bm{f}$. Such multi-level decomposition (forward transform) and reconstruction (backward transform) procedures are summarized in Algorithms 3 and 4, and illustrated in Figure 2, respectively. Here, in Algorithm 4, the notion $\text{ID}(k^{\prime})$ is defined by sorting $k^{\prime}\in\mathcal{C}_{j,k}$ increasingly to index each $k^{\prime}\in\mathcal{C}_{j,k}$ by a unique integer $\text{ID}(k^{\prime})$ in $\{1,\dots,|\mathcal{C}_{j,k}|\}$ such that $\text{ID}(k^{\prime})<\text{ID}(k^{\prime\prime})\Leftrightarrow k^{\prime}\leq k^{\prime\prime}$.
@@
(a) Forward
(b) Backward
Figure 2: (a) Forward transform: An illustration of computing coefficient vectors on a non-leaf node of $\mathscr{T}_{J}(\mathcal{G})$. The whole transform is computed iteratively in a bottom-up manner at each non-leaf node of $\mathscr{T}_{J}(\mathcal{G})$. (b) Backward Transform:An illustration of reconstruction on a non-leaf node of $\mathscr{T}_{J}(\mathcal{G})$. The whole transform is computed iteratively in a top-down manner at each non-leaf node of $\mathscr{T}_{J}(\mathcal{G})$. $\cdot]_{i,:}$ denotes the operation of extracting the $i$-th row of the matrix.
3.2 Computational Complexity
@@
Theorem 3.3.
Adopt the notations in Theorem 2.8 and Corollary 3.1. Assume that the computational complexity of CONCAT, VEC, and MAT are of order $\mathcal{O}(\text{size of input})$, and there exist positive constants $t,s,r,\alpha,m$ such that $\min\{t,s,r,\alpha,m\}\geq 1$ and
(22) $t\leq|\mathcal{C}_{j,k}|\leq s,\,s\leq t^{\alpha},\,r_{j}\leq r\leq\frac{s}{2},\,m_{j,k}\leq m$
@@
$+$ $\mathcal{O}(s\cdot r^{J-j})$ $(\text{{Matrix multiplication with }}\bm{A}^{[j,k]}),$
@@
As for Algorithm 4, it is obvious that lines 4, 5, and 6 together become the reverse of lines 5 and 6 of Algorithm 3, and that the amounts of computation are the same. Thus, the computational complexity of Algorithm 4 is the same as that of Algorithm 3.
Remark 3.5.
@@
4 Data-adaptive Graph Framelet Systems
@@
For determining the low-pass filters, we aim at learning low-pass filters so that the space $\mathbb{V}_{0}=\mathrm{span}(\bm{\Phi}_{0})$ provides the best fitting of the samples in the $l^{2}$ sense. That is, for any signal $\bm{f}\in\mathfrak{F}$, most of the energy of $\bm{f}$ is contained in its projection on $\mathbb{V}_{0}$. In such a way, we could capture the low-dimensional structure of the family $\mathfrak{F}$. Once we obtained the low-pass filters, in view of the generalized vanishing moments, i.e., orthogonality between $\mathbb{V}_{0}$ and $\mathbb{W}_{j}$, most of the coefficients in $\mathbb{W}_{j}$ are expected to be small or even zero. Therefore, in such a case, the $\mathcal{G}$-framelet system $\mathsf{GIF}_{J}(\mathcal{G})$ will provide sparse representations for all signals in $\mathfrak{F}$ (after thresholding).
The collection of matrices in $\mathbb{R}^{n\times p}$ such that $\bm{M}^{\top}\bm{M}=\bm{I}$ forms a Stiefel manifold [1] endowed with the Riemannian metric induced from the Euclidean metric of $\mathbb{R}^{n\times p}$. We denote such a manifold in $\mathbb{R}^{n\times p}$ by $\mathcal{M}^{[n,p]}$. Next, we provide the learning of the low- and high-pass filters through optimization on Stiefel manifolds.
4.1 Learning the Low-pass Filters
@@
More specifically, assuming $r_{j},j=0,\ldots,J-1$ are specified such that $r_{j}\in[\min(|\mathcal{C}_{j,1}|,\dots,|\mathcal{C}_{j,n_{j}})|-1]$, we solve the following problem for the low-pass filters:
(23) $\underset{{\left\{\bm{A}^{[j,k]}\in\mathcal{M}^{[r_{j},|\mathcal{C}_{j,k}|]}\,|\,k\in[n_{j}]\right\}_{j=0}^{J-1}}}{\arg\max}\|\bm{c}^{[0,1]}_{1}\|^{2}+\dots+\|\bm{c}^{[0,1]}_{n_{f}}\|^{2},\\$
@@
where $\bm{c}^{[0,1]}_{i}=\langle\bm{f}_{i},\bm{\Phi}_{0}\rangle$, $i\in[n_{f}]$, are the coefficients with respect to $\bm{\Phi}_{0}$ for $\bm{f}_{i}$, $i\in[n_{f}]$, computed as in Algorithm 3. Note that these coefficients only depend on the filters $\bm{A}^{[j,k]}$. By solving the problem in (23), the scaling functions $\bm{\Phi}_{0}$ and the space $\mathbb{V}_{0}$ are simultaneously determined. Moreover, the objective function in (23) is smooth on the product manifold $\mathcal{M}_{\bm{A}}:=\otimes_{j,k}\mathcal{M}^{[r_{j},|\mathcal{C}_{j,k}|]}$ formed by all $\mathcal{M}^{[r_{j},|\mathcal{C}_{j,k}|]}$. Thus, the problem above can be regarded as an optimization on Riemannian manifolds and we can choose any Riemannian optimization algorithm that is eligible to apply, e.g., see [1].
By definition (23) gives the best orthonormal vectors that best fit the training signals in $l^{2}$ sense under the constraint that they are induced by the filters $\bm{A}^{[j,k]}$. Let $\bm{M}:=\llbracket\bm{f}_{i}\rrbracket_{i\in[n_{f}]}$ denote the matrix formed by concatenating the training signals. Without requiring the orthonormal vectors are the rows of $\bm{\Phi}_{0}$, (23) becomes
$\underset{\bm{P}\in\mathbb{R}^{R_{0}\times n},\bm{P}\bm{P}^{\top}=\bm{I}}{\arg\max}\operatorname{Tr}(\bm{P}\bm{M}^{\top}\bm{M}\bm{P}^{\top}).$
This is a well-known problem which is related to principal component analysis, and one of the optimal solutions is the first $R_{0}$ right singular vectors of $\bm{M}$ (see e.g. the Eckart-Young theorem [18]). Hence, other than conducting Riemannian optimization for (23), another possible way to decide the filters $\bm{A}^{[j,k]}$ is to obtain the first $R_{0}$ right singular vectors of $\bm{M}$, then decide the entries in $\bm{A}^{[j,k]}$ such that $\Phi_{0}$ contains exactly the first $R_{0}$ right singular vectors of $\bm{M}$. However, this is guaranteed for some special cases, which needs the following lemma.
Lemma 4.1.
Let $\text{supp}(\bm{p})$ denote the support of a vector $\bm{p}$ in $\mathbb{R}^{n}$, i.e. the indices with non-zero values and $\bm{p}|_{I},I\subseteq[n]$ denote the restriction of $\bm{p}$ on a set $I$ of indices. Given vectors $\bm{p},\bm{q_{1}},\dots,\bm{q}_{r}\in\mathbb{R}^{n}$ with unit $l^{2}$ norm, if $\bigcup_{i=1}^{m}\text{supp}(\bm{q}_{i})=\text{supp}(\bm{p})$ and $\text{supp}(\bm{q}_{i})\cap\text{supp}(\bm{q}_{j})=\varnothing,i\neq j$. Then $\bm{p}=\sum_{i=1}^{m}a_{i}\bm{q}_{i}$ if and only if
$\bm{q}_{i}=\frac{\bm{p}|_{\text{supp}(\bm{q}_{i})}}{\|\bm{p}|_{\text{supp}(\bm{q}_{i})}\|},\quad a_{i}=\|\bm{p}|_{\text{supp}(\bm{q}_{i})}\|,\quad i=1,\dots,m.$
Proof 4.2.
The if part is straightforward. The only if part can be implied from the conditions $\|\bm{q}_{i}\|=1$ and $\bigcup_{i=1}^{m}\text{supp}(\bm{q}_{i})=\text{supp}(\bm{p})$ and $\text{supp}(\bm{q}_{i})\cap\text{supp}(\bm{q}_{j})=\varnothing,i\neq j$.
Then, we have the following theorem.
Theorem 4.3.
Let $\mathscr{T}_{J}({\mathcal{V}})$ be a partition tree as in Definition 2.6. Fix $r_{j}=1$ for $j=0,\dots,J-1$ in Definition 2.7. Let $\bm{p}$ be an orthonormal vector. Then, there exist filter matrices $\bm{A}^{[j,k]}\in\mathbb{R}^{r_{j}\times|\mathcal{C}_{j,k}|}$ associated with the tree node $p_{j,k}$ for $j=0,\ldots,J-1$ and $k\in[n_{j}]$, such that $\bm{\Phi}_{0}$ in the induced $\mathcal{V}$-framelet system $\mathcal{F}_{0}^{J}(\mathcal{V})$ contains exactly $\bm{p}$.
Proof 4.4.
Since $r_{j}=1$, the subspaces $\mathbb{V}^{[j,k]}$ are one-dimensional, and each $\bm{\Phi}^{[j,k]}$ can contain only one row. Note that in this case, $\mathcal{A}^{[j,k]}=\bm{A}^{[j,k]}$ and $\bm{A}^{[j,k]}$ is a row vector. Starting from $\bm{\Phi}^{[0,1]}=[\bm{p}]$, if for any $\mathcal{S}_{1,k}$, $\mathcal{S}_{1,k}\cap\mathrm{supp}(\bm{p})=\varnothing$, then let $\bm{\Phi}^{[1,k]}$ be any unit-norm row vector with support contained in $\mathcal{S}_{1,k}$, and the corresponding entry in $\bm{A}^{[j,k]}$ be zero; for the remaining $\mathcal{S}_{1,k}$ such that $\mathcal{S}_{1,k}\cap\mathrm{supp}(\bm{p})\neq\varnothing$, by (12) and Lemma 4.1, we can decide the entries of $\bm{A}^{[0,0]}$ and the matrices $\bm{\Phi}^{[1,k]}$, which correspond to the $a_{i}$ and $\bm{q}_{i}$, respectively. We can recursively perform such deductions from top to bottom until all $\bm{A}^{[j,k]}$ and $\bm{\Phi}^{[j,k]}$ are decided.
Therefore, when all the subspaces $\mathbb{V}_{j,k}$ are one-dimensional, i.e., we only seek one generalized vanishing moment, we can decide all the filters $\bm{A}^{[j,k]}$ in a top-down manner. However, if some of the subspaces $\mathbb{V}_{j,k}$ are not one-dimensional, then such a procedure fails. For example, suppose that the dimensions of $\mathbb{V}_{0}$ and all subspaces $\mathbb{V}_{1,k}$ are $2$ and that the basis of $\mathbb{V}_{0}$ is known. In this case, the filter $\bm{A}^{[0,0]}$ is essentially a row vector. By taking the first vector in the basis of $\mathbb{V}_{0}$ and applying Proposition 4.1, the filter $\bm{A}^{[0,0]}$ and the first vectors in the bases of each $\mathbb{V}_{1,k}$ are decided. However, doing the same for the second vector in the basis of $\mathbb{V}_{0}$ produces a (generally) different set of values for $\bm{A}^{[0,0]}$, which leads to a contradiction. When only $V_{0}$ is multi-dimensional, the conclusion is similar, in which Lemma 4.1 generally leads to different sets of $\bm{\Phi}^{[1,k]}$. However, we can have one fixed set. In such cases, we resort to solving (23). Such differences in learning the filters $\bm{A}^{[j,k]}$ actually present a trade-off in our systems: allowing the subspaces $\mathbb{V}_{j,k}$ to have more than one dimension gives the possibility to obtain multi-dimension $\mathbb{V}_{0}$ using fewer parameters in $\bm{A}^{[j,k]}$, with the price of losing algorithmic simplicity and optimality.
4.2 Learning the High-pass Filters